OpenStudy (anonymous):

solve for m r=2m+2z m=?

OpenStudy (zepp):

You can't solve for m but you can write m in terms of r and z.

OpenStudy (zepp):

Simply set m to one side, this is algebra, you've asked many question like this before, you can do it :)

OpenStudy (anonymous):

I really struggle with Algebra. I always have no matter how hard i work at it

OpenStudy (zepp):

We need first to set the coefficient of m and m together alone We need to have 2m alone on one side, can you do that?

OpenStudy (compassionate):

\[m = r/2 - z\]

OpenStudy (zepp):

@Compassionate You should let her to it

OpenStudy (compassionate):

That is r m = ---- - z 2

OpenStudy (compassionate):

She is lacking the mental capacity to do it. That's the problem.

OpenStudy (zepp):

That's not the problem, that's because she didn't get a good explanation.

OpenStudy (compassionate):

Give her one than.

OpenStudy (zepp):

Let's treat each monomial as one single variable; r=2m+2z Let's say r is a 2m is b 2z is c So we have \(a=b+c\)

OpenStudy (anonymous):

*then

OpenStudy (anonymous):

ok, ill figure it out. Thanks. I do have the mental capcity but my strength is in other areas and my greatest weakness is Math. People can explain till their blue in the face and it doesn't connect. I use math everyday in mywoork but its formulated for me as its business math.

OpenStudy (compassionate):

How dare you correct my grammar, @higgs

OpenStudy (anonymous):

sorry i couldn't resist the temptation

OpenStudy (compassionate):

*I Hehe.

OpenStudy (compassionate):

"Sorry, I couldn't resist the temptation." Hehe.

OpenStudy (anonymous):

I was correcting your spelling, not grammar.

OpenStudy (compassionate):

I'm concerning your grammar. :P)

OpenStudy (zepp):

As I said earlier, you must set the what's with m and itself alone on one side of an equations. 2m is b in our case, so let's set b on one side; a=b+c b=a-c Now replace all variables 2m = r - 2z We need now to set m alone, therefore we must take away the coefficient, to do that, you simply divide both side by 2\[m=\frac{r-2z}{2}\]and that could be simplified as \[m=\frac{r}{2}-\frac{2z}{2}=\frac{r}{2}-z\]

OpenStudy (anonymous):

If not, then your sentence did not make sense. Might confuse someone.

OpenStudy (compassionate):

@higgs Zepp deserves a medal.

OpenStudy (anonymous):

Yes he does.

OpenStudy (compassionate):

I love how we're spamming the hell out of this topic.

OpenStudy (anonymous):

It's closed, so it's okay. I guess.

OpenStudy (compassionate):

SPAMLALALA

OpenStudy (anonymous):

glad u all are having a blast

OpenStudy (compassionate):

This thread is full of win.